OpenStudy (anonymous):

Can someone please help with these questions? Someone who actually knows how to do this:)

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

Convert each number from scientific notation to real. 1. 2.742 x 10^4

OpenStudy (anonymous):

@learner11

OpenStudy (mrnood):

can you write 10^4 as a number? 10^1 = 10 10^2 = 100 .....

OpenStudy (anonymous):

What? I need the answer

OpenStudy (anonymous):

@imqwerty

OpenStudy (mrnood):

Oh I'm sorry I had mistaken you fro someone who actually wanted to LEARN something You asked for 'HELP' Giving oyu the answer is NOT help Try this: Can you write 10^4 as a number? 10^1 = 10 10^2 = 100 10^3 = ? 10^4 = ??

OpenStudy (mrnood):

btw I "actually know how to do this:)"

OpenStudy (mrnood):

@imqwerty why did you choose 10000? (hint - I already know - I am trying to get the poster to know!)

imqwerty (imqwerty):

sry :D i k u knw i jst wanted poster to reply :)